the perimeter of a rectangle is 46 ft. the length is 5 ft longer than the width. find the dimensions. write a system of linear equations and solve the resulting system. let x be the length and y be the width.
write the first equation
2x + 2y = ?
write the second equation.
x = y + ?
what is the length ? ft
what is the width? ? ft

Answer by Cromlix(4381)

x = length
y = width
Perimeter:
2x + 2y = 46.....1
Length is 5ft longer than width
x = y + 5......2
2x + 2y = 46....1
x - y = 5.....2
Multiply 2 by 2
2x + 2y = 46...1
2x - 2y = 10...3
Add (1) + (3)
4x = 56
x = 14
Substitute x = 14 into (2)
x - y = 5
14 - y = 5
- y = 5 - 14
- y = - 9
y = 9
x = 14
y = 9
Hope this helps
:-)